OpenStudy (anonymous):

which expression is equal to the expression below ? 3^4*5^4 ? a.(3*5) ^16 b.(3*5)^4 c.(3*5)^8 d.15^8 ? help plz thank you :D

OpenStudy (anonymous):

b.

OpenStudy (anonymous):

3^4 * 5^4 = 3*3*3*3 * 5*5*5*5 = 3*5 * 3*5 * 3*5 * 3*5 = (3*5)^4

Directrix (directrix):

3^4*5^4 = (3^1*5^1) ^4 = (3*5) ^4 --> B